OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 − 2x + 4 8x + y = 20

OpenStudy (anonymous):

i don't know how to solve this

OpenStudy (anonymous):

ok first x2-2x cancel each other out so y=+4 for the second one substitute 4 for y to get 8x+4=20 now do inverse and subtract 4 from both sides -4 -4 8x=16 /8 /8 x=2 so x=2 and y=4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@adrianperez u there

OpenStudy (anonymous):

no\[x2-2x\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@adrianperez dont forget to fan and medal those who help u